# VeriStruct: AI-assisted Automated Verification of Data-Structure Modules in Verus

### v1

@@ 1 Introduction
The power of generative AI introduces new risks to our critical digital infrastructure. AI systems have a remarkable ability to analyze code, making it much easier for a malicious actor to find and exploit vulnerabilities [1]. At the same time, more and more code is being written by or with the assistance of AI systems. While it is true that such AI systems can greatly increase the productivity of programmers, they can also introduce additional correctness errors and security vulnerabilities [2, 3]. Thus, productivity gains come at the risk of churning out mountains of buggy, insecure code.
Program verification can mitigate these risks by mathematically proving that code—whether human-written or AI-generated—is free of bugs and vulnerabilities. However, program verification relies on the addition of sophisticated logical annotations to code, including preconditions, postconditions, invariants, and auxiliary proof blocks. These annotations, which we also refer to as specifications, are used to generate verification conditions, which are then proved automatically by automated reasoning tools like satisfiability modulo theories (SMT) solvers [4]. The application of program verification in practice has been severely limited because of the enormous effort and expertise required to create such annotations.
Fortunately, the power of generative AI is also shifting the calculus on the practicality of program verification. Recently, there have been encouraging results on using large language models (LLMs) to automatically specify and verify software [5, 6, 7, 8, 9]. Our work aims to build on these efforts with the vision of one day making it possible to have AI-assisted program verification at scale and verified libraries of reusable code [10].
@@
We implement our workflow in VeriStruct, an AI-assisted automated verification framework for Verus [11, 12]. Verus is a state-of-the-art program verification extension for Rust. We target Verus because it has been successfully applied to the construction of verified systems [12], and its design allows developers to write verification annotations using a syntax closely aligned with Rust. This design not only lowers the barrier for Rust engineers to adopt formal verification but also positions Verus to have an increasing impact as Rust continues to gain traction in mainstream software development.
The input to VeriStruct consists of the Rust source code for the module to be verified and a unit test suite that illustrates the intended usage and expected behavior of the data structure. The test suite serves a dual purpose: it helps ensure that the generated formal specifications align with the developer’s intent, and it helps rule out trivial or vacuous specifications [13]. VeriStruct automatically augments the provided Rust implementation with annotations needed for program verification. This augmented code, together with the provided unit test suite, are then sent to the Verus verifier, which attempts to formally verify the assertions in the test suite. If the verification check succeeds, the formally verified code is returned to the user. Otherwise, VeriStruct attempts to extend or repair the annotations based on the feedback from the verifier.
We highlight two challenges for LLM-assisted verification of data-structure modules. The first challenge is that verifying data structures is inherently more complex than verifying a single function. Specifically, data-structure verification often requires two additional elements: (ii) a suitable mathematical abstraction that allows the verifier to reason logically about the data structure [14, 15, 16]; and ($ii$) so-called type invariants that must be preserved by operations on the data structure [8, 17]. Moreover, data-structure modules usually expose multiple methods that must be verified jointly under a shared type invariant. This stands in contrast to other recent approaches, which focus on generating a single specification artifact (e.g., proof blocks for an individual function or automated synthesis of class invariants) [7, 18, 9, 6, 8, 19, 20]. In short, our verification task extends beyond the scope of these existing techniques.
@@
To address the first challenge, we build on previous work [7] by adding two new modules: (1) a View module responsible for producing a mathematical abstraction of the data structure; and (2) a Type Invariant module, responsible for producing type invariants. For each component, we design dedicated system and user prompts that clearly articulate the task. These prompts include: (1) the objective of the component; (2) relevant background information to help the LLM better understand the task; (3) step-by-step instructions and the required output format; and (4) in-context learning examples [21]. Additionally, because not all verification components are necessary for every task, we introduce a planner module, which analyzes the verification task as an initial step and decides which components need to be generated.
To address the second challenge, we pursued two approaches. First, we developed a suite of automated repair modules to identify and correct errors in the generated annotations. Compared with the previous work [7] which focused solely on repairing proof blocks, we further introduce dedicated repair modules targeting views, type invariants, and specifications. Specifically, to facilitate the joint verification of multiple methods and tests, VeriStruct includes a repair module that refines specifications whenever the verifier reports a failed test verification. Second, we embedded structured syntax guidelines into the prompt to mitigate syntax errors in generated annotations. We sourced these guidelines from the Verus tutorial [22] and from standard Verus libraries [23].
@@
The prompt and source code of VeriStruct is available on Anonymous GitHub [24].
@@ 2 Preliminaries: Verifying a Data-Structure in Verus
A ring buffer is a fixed-size data structure that treats memory as a circular array. Ring buffers are widely used in streaming, networking, and real-time systems [25, 26]. Fig.˜1 shows the Rust implementation for a ring buffer, including the Verus annotations necessary for verifying key properties of the implementation.∎
@@ 3 The VeriStruct Workflow
The overall workflow is shown in Fig.˜2. VeriStruct builds on the general approach of the AutoVerus system [7], in that it consists of a network of modules that collaborate to complete program verification tasks. However, as noted in the challenge discussion of Sect.˜1, verifying data structures requires techniques and components not found in AutoVerus or any prior work [7, 6, 8, 9, 18].
@@ 3 Function VerifyModule(code, test)
Due to space limitations, we only present sketches of our prompts, but the full details of our prompts are publicly available [24].
@@ 4 Function GenAnnos(code, test)
Following the generated plan, the framework sequentially invokes each module $M_{i}$ (Lines 6–10). The function ExecWithSampling takes as input: (1) a module to be executed; (2) the code to be verified; and (3) the unit test suite. It executes the module, and returns the annotated code. To improve robustness, it asks for n samples during each module invocation [27]. It then selects the result yielding the maximum number of successfully verified functions. We describe the view generation module in detail below, and then describe some differentiating aspects of the other modules, which are otherwise similar.
@@
    Verus Guidelines. This part provides conceptual foundations for generating a valid View implementation. As discussed in Sect.˜1, LLMs often struggle with Verus’ syntax and verification-specific semantics. To mitigate this, our framework incorporates guidelines on Verus syntax and semantics imported from the Verus standard library documentation [23] and the official Verus tutorial [22]. For each different module in VeriStruct, a specific set of guidelines is added to the prompt. For View generation, for example, we include:
@@
    Examples [21]. This part provides examples of View implementations for other verified data structures (e.g., doubly-linked lists). These examples are kept the same across (and are disjoint from) all benchmarks.
@@ 6 Evaluation
We evaluate VeriStruct on a benchmark set comprising eleven Rust data structure modules drawn from the Verus GitHub repository [11] as well as other open-source repositories. A list of the benchmarks is shown in Table˜1. The benchmarks cover a set of commonly-used data structures, including a ring buffer, a vector implementation with set abstractions and common algorithms, a polymorphic option type, tree-based modules for standalone nodes and a treemap, a bitmap with 64-bit blocks, and concurrent data structures such as a read–write lock and an invariant framework for shared state. Each benchmark contains 5–21 functions to verify, where the number of functions refers to the total number of member methods in the data structure module and the corresponding test functions. In total, our benchmark set includes 129 functions to be verified across all modules.
@@
We run VeriStruct and the baseline on every benchmark individually. For each benchmark we first execute VeriStruct to completion and then record how many LLM invocations it required. We observe that the maximum number of LLM invocations for VeriStruct is 13, so we run the baseline on each benchmark with 13 LLM invocations. All experiments are conducted on a server running Ubuntu 22.04.5 LTS with a 24-core Intel(R) Core(TM) i9-12900K CPU and 64 GB RAM. We use OpenAI o1 [28] as the back-end model for both VeriStruct and the baseline. We instantiate VeriStruct with n=3 and m=5; that is, the number of samples generated per module is 3, and the maximum number of repair rounds is set to 5.
@@ Results
We summarize the results in Tables˜2 and 3, VeriStruct successfully solves 10 out of 11 benchmarks. For the only unsolved benchmark (Node), VeriStruct still verifies 11 out of 12 functions. In contrast, the baseline solves only 4 out of 11 benchmarks. In terms of the total number of functions verified, VeriStruct verifies 128 out of 129 (99.2%) functions, whereas the baseline verifies only 52. In summary, VeriStruct solves substantially more benchmarks and verifies significantly more functions, demonstrating that our approach improves the capabilities of LLMs to generate correct annotations.
@@ 7 Related Work
Towards AI-assisted automated program verification, there is a long research line focusing on verifying the correctness of either an individual function [6, 7, 9, 20, 19] or the type invariant for a given module [8]. In contrast, this paper targets the verification of data structure modules, which requires the synthesis of mathematical abstractions, structural invariants, and specifications and proof blocks spanning multiple methods. Consequently, none of the previous work can be applied in our setting.
Traditional formal-methods tools (e.g., Frama-C [29], Why3 [30], Dafny [31], Coq [32], and Isabelle [33]) provide strong correctness guarantees but require developers to manually construct detailed specifications and proofs. Thus, they impose substantial annotation overhead and demand significant expertise. In contrast, VeriStruct leverages large language models to automatically synthesize annotations. As a result, VeriStruct bridges the gap between fully manual formal verification and prior LLM-based assistants, achieving greater automation without sacrificing rigor.
@@ 8 Conclusion and Future Work
We conclude with several promising directions for future work. A natural next step is to extend VeriStruct to support more complex verification tasks. First, we could integrate retrieval-augmented generation (RAG) [34], which could help identify useful lemmas from the Verus standard library [35], strengthening proof synthesis. Second, we could additionally support the synthesis of Verus’ resource algebra library [12] for verifying complex concurrent data structures. Finally, applying constrained decoding techniques [36] could alleviate syntax errors in generated annotations.
@@
At present, our approach requires users to manually provide a comprehensive unit test suite. However, constructing such suites, especially those that achieve high coverage and include corner cases, can be both time-consuming and error-prone. Recent studies have shown that large language models are highly capable of generating high-quality test cases, including those targeting rare or edge conditions [37, 38]. Building on these advances, we plan to incorporate LLM-based unit test generation into our framework, thereby further reducing manual effort and enhancing the completeness of the verification process.

### v2

@@ 1 Introduction
The power of generative AI introduces new risks to our critical digital infrastructure. AI systems have a remarkable ability to analyze code, making it much easier for a malicious actor to find and exploit vulnerabilities [BBB+23]. At the same time, more and more code is being written by or with the assistance of AI systems. While it is true that such AI systems can greatly increase the productivity of programmers, they can also introduce additional correctness errors and security vulnerabilities [insecure1, bug1]. Thus, productivity gains come at the risk of churning out mountains of buggy, insecure code.
Program verification can mitigate these risks by mathematically proving that code—whether human-written or AI-generated—is free of bugs and vulnerabilities. However, program verification relies on the addition of sophisticated logical annotations to code, including preconditions, postconditions, invariants, and auxiliary proof blocks. These annotations, which we also refer to as specifications, are used to generate verification conditions, which are then proved automatically by automated reasoning tools like satisfiability modulo theories (SMT) solvers [BSST21]. The application of program verification in practice has been severely limited because of the enormous effort and expertise required to create such annotations.
Fortunately, the power of generative AI is also shifting the calculus on the practicality of program verification. Recently, there have been encouraging results on using large language models (LLMs) to automatically specify and verify software [clover, cav24llm, autoverus, classinvgen, dafnysinglemethod]. Our work aims to build on these efforts with the vision of one day making it possible to have AI-assisted program verification at scale and verified libraries of reusable code [verilib].
@@
We implement our workflow in VeriStruct, an AI-assisted automated verification framework for Verus [verus, verus-practical]. Verus is a state-of-the-art program verification extension for Rust. We target Verus because it has been successfully applied to the construction of verified systems [verus-practical], and its design allows developers to write verification annotations using a syntax closely aligned with Rust. This design not only lowers the barrier for Rust engineers to adopt formal verification but also positions Verus to have an increasing impact as Rust continues to gain traction in mainstream software development.
The input to VeriStruct consists of the Rust source code for the module to be verified and a unit test suite that illustrates the intended usage and expected behavior of the data structure. The test suite serves a dual purpose: it helps ensure that the generated formal specifications align with the developer’s intent, and it helps rule out trivial or vacuous specifications [test-intent]. VeriStruct automatically augments the provided Rust implementation with annotations needed for program verification. This augmented code, together with the provided unit test suite, are then sent to the Verus verifier, which attempts to formally verify the assertions in the test suite. If the verification check succeeds, the formally verified code is returned to the user. Otherwise, VeriStruct attempts to extend or repair the annotations based on the feedback from the verifier.
We highlight two challenges for LLM-assisted verification of data-structure modules. The first challenge is that verifying data structures is inherently more complex than verifying a single function. Specifically, data-structure verification often requires two additional elements: (ii) a suitable mathematical abstraction that allows the verifier to reason logically about the data structure [reynolds-seplogic, parkinson-bierman, ohearn-cacm]; and ($ii$) so-called type invariants that must be preserved by operations on the data structure [classinvgen, pldi-invariant]. Moreover, data-structure modules usually expose multiple methods that must be verified jointly under a shared type invariant. This stands in contrast to other recent approaches, which focus on generating a single specification artifact (e.g., proof blocks for an individual function or automated synthesis of class invariants) [autoverus, dafny-annotator, dafnysinglemethod, cav24llm, classinvgen, lemur, loopy]. In short, our verification task extends beyond the scope of these existing techniques.
@@
To address the first challenge, we build on previous work [autoverus] by adding two new modules: (1) a View module responsible for producing a mathematical abstraction of the data structure; and (2) a Type Invariant module, responsible for producing type invariants. For each component, we design dedicated system and user prompts that clearly articulate the task. These prompts include: (1) the objective of the component; (2) relevant background information to help the LLM better understand the task; (3) step-by-step instructions and the required output format; and (4) in-context learning examples [incontextlearning]. Additionally, because not all verification components are necessary for every task, we introduce a planner module, which analyzes the verification task as an initial step and decides which components need to be generated.
To address the second challenge, we pursued two approaches. First, we developed a suite of automated repair modules to identify and correct errors in the generated annotations. Compared with the previous work [autoverus] which focused solely on repairing proof blocks, we further introduce dedicated repair modules targeting views, type invariants, and specifications. Specifically, to facilitate the joint verification of multiple methods and tests, VeriStruct includes a repair module that refines specifications whenever the verifier reports a failed test verification. Second, we embedded structured syntax guidelines into the prompt to mitigate syntax errors in generated annotations. We sourced these guidelines from the Verus tutorial [verus-tutorial] and from standard Verus libraries [verus-stdlib].
@@
The prompt and source code of VeriStruct is available on GitHub at https://github.com/ChuyueSun/VeriStruct[artifact]
@@ 2 Preliminaries: Verifying a Data-Structure in Verus
A ring buffer is a fixed-size data structure that treats memory as a circular array. Ring buffers are widely used in streaming, networking, and real-time systems [tock, knuth1997art]. Fig.˜1 shows the Rust implementation for a ring buffer, including the Verus annotations necessary for verifying key properties of the implementation.∎
@@ 3 The VeriStruct Workflow
The overall workflow is shown in Fig.˜2. VeriStruct builds on the general approach of the AutoVerus system [autoverus], in that it consists of a network of modules that collaborate to complete program verification tasks. However, as noted in the challenge discussion of Sect.˜1, verifying data structures requires techniques and components not found in AutoVerus or any prior work [autoverus, cav24llm, classinvgen, dafnysinglemethod, dafny-annotator].
@@ 3 Function VerifyModule(code, test)
Due to space limitations, we only present sketches of our prompts, but the full details of our prompts are publicly available [artifact].
@@ 4 Function GenAnnos(code, test)
Following the generated plan, the framework sequentially invokes each module $M_{i}$ (Lines 6–10). The function ExecWithSampling takes as input: (1) a module to be executed; (2) the code to be verified; and (3) the unit test suite. It executes the module, and returns the annotated code. To improve robustness, it asks for n samples during each module invocation [sampling]. It then selects the result yielding the maximum number of successfully verified functions. We describe the view generation module in detail below, and then describe some differentiating aspects of the other modules, which are otherwise similar.
@@
    Verus Guidelines. This part provides conceptual foundations for generating a valid View implementation. As discussed in Sect.˜1, LLMs often struggle with Verus’ syntax and verification-specific semantics. To mitigate this, our framework incorporates guidelines on Verus syntax and semantics imported from the Verus standard library documentation [verus-stdlib] and the official Verus tutorial [verus-tutorial]. For each different module in VeriStruct, a specific set of guidelines is added to the prompt. For View generation, for example, we include:
@@
    Examples [incontextlearning]. This part provides examples of View implementations for other verified data structures (e.g., doubly-linked lists). These examples are kept the same across (and are disjoint from) all benchmarks.
@@ 6 Evaluation
We evaluate VeriStruct on a benchmark set comprising eleven Rust data structure modules drawn from the Verus GitHub repository [verus] as well as other open-source repositories. A list of the benchmarks is shown in Table˜1. The benchmarks cover a set of commonly-used data structures, including a ring buffer, a vector implementation with set abstractions and common algorithms, a polymorphic option type, tree-based modules for standalone nodes and a treemap, a bitmap with 64-bit blocks, and concurrent data structures such as a read–write lock and an invariant framework for shared state. Each benchmark contains 5–21 functions to verify, where the number of functions refers to the total number of member methods in the data structure module and the corresponding test functions. In total, our benchmark set includes 129 functions to be verified across all modules.
@@
We run VeriStruct and the baseline on every benchmark individually. For each benchmark we first execute VeriStruct to completion and then record how many LLM invocations it required. We observe that the maximum number of LLM invocations for VeriStruct is 13, so we run the baseline on each benchmark with 13 LLM invocations. All experiments are conducted on a server running Ubuntu 22.04.5 LTS with a 24-core Intel(R) Core(TM) i9-12900K CPU and 64 GB RAM. We use OpenAI o1 [o1] as the back-end model for both VeriStruct and the baseline. We instantiate VeriStruct with n=3 and m=5; that is, the number of samples generated per module is 3, and the maximum number of repair rounds is set to 5.
@@ Results
We summarize the results in Tables˜2 and 5, VeriStruct successfully solves 10 out of 11 benchmarks. For the only unsolved benchmark (Node), VeriStruct still verifies 11 out of 12 functions. In contrast, the baseline solves only 4 out of 11 benchmarks. In terms of the total number of functions verified, VeriStruct verifies 128 out of 129 (99.2%) functions, whereas the baseline verifies only 52. In summary, VeriStruct solves substantially more benchmarks and verifies significantly more functions, demonstrating that our approach improves the capabilities of LLMs to generate correct annotations.
@@
6.1 Robustness Evaluation
We repeat the eleven-task evaluation five times to quantify robustness. Each run executes the full workflow with a fresh cache, identical sampling parameters, and up to five repair rounds per benchmark, giving 55 independent executions that isolate variance due to stochastic sampling.
Per-Benchmark Outcomes and Aggregate Statistics
Nine benchmarks verify in every run, Bitmap succeeds in three of five runs, and RingBuffer succeeds once. Aggregating those outcomes yields an 89.1% success rate (49/55 executions) with a 3.6% standard deviation across runs. Table˜3 reports that benchmarks consume 21,921 tokens on average with a 2.7% coefficient of variation in the total per-run token budget, while Table˜4 shows the 10,649–38,514 token range for individual benchmarks.
Table 3: Aggregate robustness statistics across the five runs using token counts.
Statistic Success Rate Avg. Tokens / Benchmark Avg. Input Avg. Output
Mean 89.1% 21,921 8,439 13,481
Std. Dev. 3.6% 587 151 512
Coefficient of Variation 2.7% 1.8% 3.8%
Table 4: Token efficiency by benchmark, grouped by reliability tier. “Token/Success” divides the average total tokens by the empirical success probability.
Benchmark Success Rate Avg. Tokens Token/Success
Tier 1: Highly Reliable (100% success)
Transfer 5/5 (100%) 10,649 10,649
Invariants 5/5 (100%) 12,414 12,414
RwLockVstd 5/5 (100%) 12,846 12,846
Option 5/5 (100%) 17,482 17,482
Atomics 5/5 (100%) 19,326 19,326
Node 5/5 (100%) 33,885 33,885
Treemap 5/5 (100%) 38,514 38,514
Vectors 5/5 (100%) 24,394 24,394
SetFromVec 5/5 (100%) 16,368 16,368
Tier 2: Variable Performance (60% success)
Bitmap 3/5 (60%) 23,324 38,873
Tier 3: Challenging (20% success)
RingBuffer 1/5 (20%) 27,250 136,250
Reliability Tiers and Insights
The benchmarks fall into three tiers: nine highly reliable modules that cover core data structures and concurrency primitives, one variable case (Bitmap) that succeeds 60% of the time, and one challenging case (RingBuffer) that rarely converges. Tier 1 problems verify with 10k–39k tokens each, Tier 2 requires roughly 39k tokens per success, and Tier 3 exceeds 136k tokens per success (Table˜4). Because per-run totals vary by under 3% (Table˜3), these token-efficiency gaps mainly reflect inherent difficulty rather than sampling noise, clarifying which cases demand new reasoning techniques beyond repeated attempts.
@@ 7 Related Work
Towards AI-assisted automated program verification, there is a long research line focusing on verifying the correctness of either an individual function [cav24llm, autoverus, dafnysinglemethod, loopy, lemur] or the type invariant for a given module [classinvgen]. In contrast, this paper targets the verification of data structure modules, which requires the synthesis of mathematical abstractions, structural invariants, and specifications and proof blocks spanning multiple methods. Consequently, none of the previous work can be applied in our setting.
Traditional formal-methods tools (e.g., Frama-C [framac], Why3 [why3], Dafny [dafny], Coq [coq], and Isabelle [isabelle]) provide strong correctness guarantees but require developers to manually construct detailed specifications and proofs. Thus, they impose substantial annotation overhead and demand significant expertise. In contrast, VeriStruct leverages large language models to automatically synthesize annotations. As a result, VeriStruct bridges the gap between fully manual formal verification and prior LLM-based assistants, achieving greater automation without sacrificing rigor.
@@ 8 Conclusion and Future Work
We conclude with several promising directions for future work. A natural next step is to extend VeriStruct to support more complex verification tasks. First, we could integrate retrieval-augmented generation (RAG) [RAG], which could help identify useful lemmas from the Verus standard library [vstd], strengthening proof synthesis. Second, we could additionally support the synthesis of Verus’ resource algebra library [verus-practical] for verifying complex concurrent data structures. Finally, applying constrained decoding techniques [constrained-decoding] could alleviate syntax errors in generated annotations.
@@
At present, our approach requires users to manually provide a comprehensive unit test suite. However, constructing such suites, especially those that achieve high coverage and include corner cases, can be both time-consuming and error-prone. Recent studies have shown that large language models are highly capable of generating high-quality test cases, including those targeting rare or edge conditions [highcovtest, liu2025llmpoweredtestcasegeneration]. Building on these advances, we plan to incorporate LLM-based unit test generation into our framework, thereby further reducing manual effort and enhancing the completeness of the verification process.